Gene purchased all of the stock of Parkside Corporation for $180,000. During 2015, the corporation redeemed one-third of the stock for $270,000. It has been determined that the transaction is properly classified as an IRC Section 301 redemption transaction. There have not been any intervening transactions that impact Genes basis in the stock of Parkside. What is Genes basis in the remaining stock of Parkside? A. $0, B. $60,000, C, $120,000, D. $180,000
Answer #1

Total cost of shares = 180000

Share cost which is sold = 180000 * 1/3 = 60000

Remaining stock = 180000 - 60000

Remaining stock value = 120000

Option C is correct

Please please upvote it at the end thank you so much